OpenStudy (camerondoherty):

How to solve \[2x^2 -10x + 3 = 0\] by completing the square?

OpenStudy (chosenmatt):

what do you think so far?

OpenStudy (camerondoherty):

Well, I was thinking to first divide both sides by 2, so you can have x^2 instead of 2x^2

OpenStudy (anonymous):

use the quadratic formula instead, it is easier in this case

OpenStudy (chosenmatt):

^^

OpenStudy (chosenmatt):

yeah its easier

OpenStudy (camerondoherty):

I would, but the question asks to complete the square cx

OpenStudy (anonymous):

it makes no difference the formula is identical to completing the square

OpenStudy (anonymous):

just does it as a formula instead of in steps

OpenStudy (chosenmatt):

do it as a foruma which is what @satellite73 said im gonna go help some other people looks like you got all the help you need @camerondoherty

OpenStudy (anonymous):

\[2x^2-10x=3\] is a start, then divide by 2 \[x^2-5x=\frac{3}{2}\] and now with annoying fractions go right to \[(x-\frac{5}{2})^2=\frac{3}{2}+\frac{9}{4}\]

OpenStudy (solomonzelman):

\(\large\color{blue}{ 2x^2-10+3=0 }\) \(\large\color{blue}{ 2x^2-10+3\color{red}{-3}=0\color{red}{-3} }\) \(\large\color{blue}{ 2x^2-10=-3 }\) \(\large\color{blue}{ 2(x^2-5)=-3 }\) \(\large\color{blue}{ 2(x^2-5\color{green}{+\frac{\LARGE 5^2}{\LARGE 2^2}}\color{red}{-\frac{\LARGE 5^2}{\LARGE 2^2}})=-3 }\) \(\large\color{blue}{ 2(x^2-5\color{green}{+\frac{\LARGE 5^2}{\LARGE 2^2}})-2\color{red}{\frac{\LARGE 5^2}{\LARGE 2^2}}=-3 }\)

OpenStudy (anonymous):

2x^2 -10x + 3 = 0 2x2−10x+3=0 Step 1: Use quadratic formula with a=2, b=-10, c=3. x= −b±sqrt(b2−4ac) 2a x= −(−10)±sqrt((−10)2−4(2)(3)) 2(2) x= 10±sqrt(76) 4 x=4.6794494717703365,0.320550528229663 Answer: x=4.6794494717703365,0.320550528229663

OpenStudy (solomonzelman):

2 many, lol

OpenStudy (anonymous):

oh whoops it is \[2x^2-10x+3=0\] so really \[(x-\frac{5}{2})^2=-\frac{3}{2}+\frac{9}{4}\]

OpenStudy (anonymous):

solution is x=1/2(5-+sqr19

OpenStudy (camerondoherty):

Hm, so \[(x-\frac{5}{2})^2=-\frac{3}{2}+\frac{9}{4}\] Adding the fractions on the side \[(x-\frac{5}{2})^2=\frac{3}{4}\] Then solving the other side \[x^2-5x+\frac{ 25 }{ 4 }= \frac{ 3 }{ 4 }\] I'm stuck here.

OpenStudy (solomonzelman):

perfect square trinomial

OpenStudy (solomonzelman):

the left side

OpenStudy (camerondoherty):

\[x^2-5x+\frac{ 22 }{ 4 }= 0\]

OpenStudy (solomonzelman):

what you had before is correct. \(\large\color{blue}{(a+b)^2=a^2+2ab+b^2 }\) \(\large\color{blue}{(a+b)^2=x^2+2(x)(5/2)+(5/2)^2 }\)

OpenStudy (camerondoherty):

Oh left side. Got left/right confused, lol

OpenStudy (solomonzelman):

See, compare it to that

OpenStudy (solomonzelman):

\(\large\color{blue}{x^2+5x+25/4=3/4 }\) is really, \(\large\color{blue}{x^2+5(x)(5/2)+(5/2)^2=3/4 }\)

OpenStudy (solomonzelman):

\(\large\color{blue}{a^2+2(a)(b)+(b)^2=>~(a+b)^2 }\)

OpenStudy (freckles):

\[ax^2+bx+c=0 \\ \text{ first step for your problem: subtract } c \text{ on both sides } \\ ax^2+bx=-c \\ \text{ second step for your problem: factor } a \text{ from both terms on left side } \\ \text{ but don't freak out because there isn't an } a \text{ next to the } bx \\ \text{ because you can put an } a \text{ there as long as you also divide by } a \text{ next to it also } \\ ax^2+\frac{a}{a}bx=-c \\ \text{ so factor out } a \text{ from both terms on the left } \\ a(x^2+\frac{1}{a}bx)=-c \\ a(x^2+\frac{b}{a}x)=-c\] \[\text{ third step: inside the parenthesis on the left side } \text{ add in a number that will complete the square } \\ \text{ but whatever you do to one side do to the other } \] \[a(x^2+\frac{b}{a}x+(\frac{b}{2a})^2)=-c+a(\frac{b}{2a})^2\] \[\text{ notice inside the ( ) I just added in the square of the number next to } x \text{ divided by 2 }\] \[\text{ but that number I added in was being multiplied by } a \] \[\text{ So I actually added on the left } a(\frac{b}{2a})^2\] \[\text{ So I also need to add that number same number on the right }\] So that is what I have above: \[a(x^2+\frac{b}{a}x+(\frac{b}{2a})^2)=-c+a(\frac{b}{2a})^2\] \[\text{ Now the whole reason I did this was so I could } \\ \text{ write that number on the left next to the number } a \text{ as something squared }\] \[\text{ so we have: } a(x+\frac{b}{2a})^2=-c+a(\frac{b}{2a})^2\] \[\text{ now I'm going to divide both sides by } a \] \[(x+\frac{b}{2a})^2=\frac{-c}{a}+\frac{a}{a}(\frac{b}{2a})^2\] \[(x+\frac{b}{2a})^2=\frac{-c}{a}+(\frac{b}{2a})^2\] I'm going to go ahead and write the right hand side as one fraction: \[(x+\frac{b}{2a})^2=\frac{-c}{a}+\frac{b^2}{4a^2}\] \[(x+\frac{b}{2a})^2=\frac{-c(4a)}{a(4a)}+\frac{b^2}{4a^2}\] \[(x+\frac{b}{2a})^2=\frac{-4ac+b^2}{4a^2}\] \[(x+\frac{b}{2a})^2=\frac{b^2-4ac}{4a^2}\] Now recall the objective is to isolate x. But oh no it is inside some square thingy... So let's take the square root of both sides : \[(x+\frac{b}{2a})=\pm \sqrt{ \frac{b^2-4ac}{4a^2}}\] \[x+\frac{b}{2a}=\pm \sqrt{ \frac{b^2-4ac}{4a^2}}\] \[x+\frac{b}{2a}= \pm \frac{\sqrt{b^2-4ac}}{\sqrt{4a^2}}\] \[x+\frac{b}{2a}= \pm \frac{\sqrt{b^2-4ac}}{\pm 2 a}\] We don't need plus or minus twice: \[x+\frac{b}{2a}= \pm \frac{\sqrt{b^2-4ac}}{ 2a} \] Subtracting b/2a on both sides give: \[x=\frac{-b}{2a} \pm \frac{\sqrt{b^2-4ac}}{2a}\] And you can combine the fractions here and thus the quadratic formula is born. \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] Now the reason I wrote this... Is because if you can understand how to complete the square with just "letters" involved then you can do it for any numbers these "letters" can represent. \[\text{ Also I hope you notice } a \text{ cannot be zero }\] :)

OpenStudy (solomonzelman):

the old equation editor, but very nice!

OpenStudy (solomonzelman):

`\(\large\color{slate}{` .... `}\)`

OpenStudy (solomonzelman):

anyways, back to math

OpenStudy (acxbox22):

wow you just proved the quadratic equation @freckles

OpenStudy (solomonzelman):

we need completing the sq/ tho

OpenStudy (solomonzelman):

well, not a lover of abstract math...

OpenStudy (freckles):

Well the method I show to get the quadratic formula involves completing the square... If you can understand all of those steps I took with "letters" then he can do it with the numbers those "letters" can represent.

OpenStudy (camerondoherty):

Sorry, The equation button went below and I couldnt reach it I have to rewrite all of this cx In my opinion, that may as well be a tutorial, its great enough to be c: Thank you for writing it *u* It really helps me understand! So, Im not going to write in in latex, but I'm going to solve it on paper c: Last time I tried writing it in latex, I lost it all, so give me a sec...

OpenStudy (freckles):

Retyped because some of the lines got cut off. \[ax^2+bx+c=0 \\ \text{ first step for your problem: subtract } c \text{ on both sides } \\ ax^2+bx=-c \\ \text{ second step for your problem: factor } a \text{ from both terms on left side } \\ \text{ but don't freak out because there isn't an } a \text{ next to the } bx \\ \text{ because you can put an } a \text{ there as long as you also divide by } a \\ \text{ next to it also } \\ ax^2+\frac{a}{a}bx=-c \\ \text{ so factor out } a \text{ from both terms on the left } \\ a(x^2+\frac{1}{a}bx)=-c \\ a(x^2+\frac{b}{a}x)=-c \\ \text{ third step: inside the parenthesis on the left side } \\ \text{ add in a number that will } \\ \text{ complete the square but whatever you do to one side do to the other } \\ a(x^2+\frac{b}{a}x+(\frac{b}{2a})^2)=-c+a(\frac{b}{2a})^2 \\ \text{ notice inside the ( ) I just added in the square of the number next to } \\ x \text{ divided by 2 } \\ \text{ but that number I added in was being multiplied by } a \\ \text{ So I actually added on the left } a(\frac{b}{2a})^2 \\ \text{ so I need to add } a(\frac{b}{2a})^2 \text{ on the right as well } \\ \text{ So that is what I have above: } \\ a(x^2+\frac{b}{a}x+(\frac{b}{2a})^2)=-c+a(\frac{b}{2a})^2 \\ \text{ Now the whole reason I did this was so I could } \\ \text{ write that number on the left next to the number } a \text{ as something squared } \\ \text{ so we have: } a(x+\frac{b}{2a})^2=-c+a(\frac{b}{2a})^2 \\ \text{ now I'm going to divide both sides by } a \\ (x+\frac{b}{2a})^2=\frac{-c}{a}+\frac{a}{a}(\frac{b}{2a})^2 \\ (x+\frac{b}{2a})^2=\frac{-c}{a}+(\frac{b}{2a})^2 \\ \text{ I'm going to go ahead and write the right hand side as one fraction: } \\ (x+\frac{b}{2a})^2=\frac{-c}{a}+\frac{b^2}{4a^2} \\ (x+\frac{b}{2a})^2=\frac{-c(4a)}{a(4a)}+\frac{b^2}{4a^2} \\ (x+\frac{b}{2a})^2=\frac{-4ac+b^2}{4a^2} \\ (x+\frac{b}{2a})^2=\frac{b^2-4ac}{4a^2} \\ \text{ Now recall the objective is to isolate x. } \\ \text{ But oh no it is inside some square thingy... } \\ \text{ So let's take the square root of both sides : } \\ (x+\frac{b}{2a})=\pm \sqrt{ \frac{b^2-4ac}{4a^2}} \\ x+\frac{b}{2a}=\pm \sqrt{ \frac{b^2-4ac}{4a^2}} \\ x+\frac{b}{2a}= \pm \frac{\sqrt{b^2-4ac}}{\sqrt{4a^2}} \\ x+\frac{b}{2a}= \pm \frac{\sqrt{b^2-4ac}}{ 2a} \\ \text{ Subtracting } \frac{ b}{2a } \text{ on both sides give: } \\ x=\frac{-b}{2a} \pm \frac{\sqrt{b^2-4ac}}{2a} \\ \text{ And you can combine the fractions here and thus } \\ \text{ the quadratic formula is born. }\\ x=\frac{-b \pm \sqrt{b^2-4ac}}{2a} \\ \text{ Now the reason I wrote this... } \\ \text{ is because if you can understand how to } \\ \text{ complete the square with just "letters" involved } \\ \text{ then you can do it for any numbers these "letters" can represent. } \\ \text{ Also I hope you notice } a \text{ cannot be zero } \]

OpenStudy (camerondoherty):

You re-wrote it? I kinda feel bad now cx Thank You so much... Omg I don't know how I can thank you enough You showed it just so thouroghly and so easy to understand!! c:

OpenStudy (camerondoherty):

When I solved by plugging in the values, I got \[\begin{array}{*{20}c} {x = \frac{{ 5 \pm \sqrt {3} }}{{2}}} & \\ \end{array}\]

OpenStudy (freckles):

\[2x^2 -10x + 3 = 0 \\ 2x^2-10x=-3 \\ 2(x^2-5x)=-3 \\ 2(x^2-5x+(\frac{5}{2})^2)=-3+2(\frac{5}{2})^2 \\ \text{ now remember we actually added in } 2 \cdot (\frac{5}{2})^2 \\ \text{ on the left } \\ \text{ so we also needed to have added } 2 \cdot (\frac{5}{2})^2 \\ \text{ so on the left next to the 2 we can write that as something squared now } \\ 2(x-\frac{5}{2})^2=-3+2 \cdot (\frac{5}{2})^2\]

OpenStudy (freckles):

I think we are suppose to get something different in the square root

OpenStudy (freckles):

like i think instead of 3 you should have something else there

OpenStudy (camerondoherty):

So, I did something wrong then cx Let me do the math again, one sec cx

OpenStudy (freckles):

ok

OpenStudy (acxbox22):

you can always use quadratic formula to check \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2 }\] \[x \frac{ 10\pm \sqrt{100-4(2)(3)} }{ 4 }\]

OpenStudy (camerondoherty):

I multiplied wrong, cx It's supposed to be: \[\begin{array}{*{20}c} {x = \frac{{ 5 \pm \sqrt 19} }}{{2a}}} & \\ \end{array}\]

OpenStudy (camerondoherty):

Wait, er, Broken Latex \[\begin{array}{*{20}c} {x = \frac{{ - b \pm \sqrt {19} }}{{2a}}} \\ \end{array}\]

OpenStudy (camerondoherty):

Wait, no cx \[\begin{array}{*{20}c} {x = \frac{{ 5 \pm \sqrt {19} }}{{2}}} \\ \end{array}\]

OpenStudy (acxbox22):

that looks right to me

OpenStudy (freckles):

\[2x^2 -10x + 3 = 0 \\ 2x^2-10x=-3 \\ 2(x^2-5x)=-3 \\ 2(x^2-5x+(\frac{5}{2})^2)=-3+2(\frac{5}{2})^2 \\ \text{ now remember we actually added in } 2 \cdot (\frac{5}{2})^2 \\ \text{ on the left } \\ \text{ so we also needed to have added } 2 \cdot (\frac{5}{2})^2 \\ \text{ so on the left next to the 2 we can write that as something squared now } \\ 2(x-\frac{5}{2})^2=-3+2 \cdot (\frac{5}{2})^2 \\ 2(x-\frac{5}{2})^2=-3+2 \cdot \frac{25}{4} \\ \text{ divide both sides by 2 } \\ \frac{2}{2}(x-\frac{5}{2})^2=\frac{-3}{2}+ \frac{2}{2} \cdot \frac{25}{4} \\ (x-\frac{5}{2})^2=\frac{-3}{2}+\frac{25}{4} \] then combine fractions on left then take square root of both sides

OpenStudy (freckles):

combine fractions on right*

OpenStudy (freckles):

yes

OpenStudy (acxbox22):

from the quadratic formula